A bus starting from rest moves with a uniform acceleration of 0.1 m s-2 for 2 minutes. Find (a) the speed acquired, (b) the distance travelled.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


Initial speed of the bus, u= 0 m/s
Acceleration, a = 0.1 m/s^2
Time taken, t = 2 minutes = 120 s
(a) v= u + at
v= 0 + 0×1 × 120
v= 12 ms-1
(b) According to the third equation of motion, v^2- u^2= 2as
s is the distance covered by the bus
(12)^2- (0)^2= 2(0.1) s
s = 720 m
Speed acquired finally by the bus is 12 m/s.
Distance travelled by the bus is 720 m.




Thanks